OpenStudy (anonymous):

Medals ANd Fan Given What is the simplified solution to five over eighteen divided by 5 over twelve? one over three two over three seventeen over twenty-three twenty-five over two hundred sixteen

OpenStudy (anonymous):

Do you know how to multiply fractions? Because dividing them is the same as multiplying the recipical.

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

So: \[\frac{ 5 }{ 18 }\div \frac{ 5 }{ 12 }\]

OpenStudy (anonymous):

Is the same thing as: \[\frac{ 5 }{ 18 }\div \frac{ 12 }{ 5 }\]

OpenStudy (anonymous):

Oops, I meant times on that last one ^^

OpenStudy (anonymous):

Then you cross multiply